OpenStudy (lolololhelpme):

Suppose Y varies inversely with x. If x=7 when y=-2, find the value of x when y=-6

OpenStudy (wckeller):

\[y=\frac{ ? }{ x } \] \[-2=\frac{ ? }{ 7 }\] -2*7 = -14 ?=-14 \[-6=\frac{ -14 }{ x }\] \[-6x=-14\] \[x=\frac{ -14 }{ -6 }\] Reduced=\[x=\frac{ 7 }{ 3 }\] Now I have to say I'm not entirely sure if I did this correctly, but it's the best I could come up with.